<<TableOfContents(3)>>

= Agenda/Minutes =
[[https://zoom.us/j/203743987|Zoom Meeting]]

----------
= CRY+VIS Interface Meeting =
Date: 2023/7/28 13:00-13:15

Participants: Takahashi, Uchiyama, Yokozawa, Ushiba, Aso, Miyoki, Ikeda, Akutsu, Sato

[[https://paper.dropbox.com/folder/show/AcceptanceTest-e.1gg8YzoPEhbTkrhvQwJ2zzo3dHVh5sJGieDgbm2n3h60mGqotBeE|Acceptance check document]]

== Suspensions ==
 * Type-A: 
 * Type-B: 
 * Type-Bp: 
 * Type-C: 

== Other items ==
 * Oplev: 
 * ACC: 
 * Commissioning: High power operation. OMC PD became noisy,
 * Vacuum system: Leak test of the OMC chamber. Leak at Y-arm was due to the CC-10.
 * Heater control: 
 * TMS: 

== Post O4a ==
 * Measurements of the DAC noise: The next time on 8/4.
 * Measurements on the sensor correction in SR2: 1 night.
 * Optimization of the damping control in PRs: Measurements 2 days, Test 3 days.
 * IP control with the ACCs: Postponed the long-term test in IY.
 * In-vacuum accelerometer for IOO: 
 * Hight adjustments of Type-A in the cooling: [[https://paper.dropbox.com/doc/Hight-adjustment-in-cooling--B7ii3omB81JIm6khnBt_7fcaAQ-swWQ5h6M6lSs8yLp5D6dA|The procedure]]. Cooling was started on 7/27.
 * Noise budget of the suspensions: 
 * Temperature control for the Oplev: Devices of the heater controller are ready.

== Discussion ==